# Towards AI-Architecture Liberty: A Comprehensive Survey on Designing and Collaborating Virtual Architecture by Deep Learning in the Metaverse

### v3

@@ 1.3. Scope and Structure
While the intersection of deep learning and architecture encompasses various aspects, including design cognition (e.g., feature extraction, image recognition, design evaluation), design generation (e.g., floor plans, rendering, form generation), and assisting tools (e.g., efficiency optimization, strategy prediction), our investigation focuses specifically on articles where deep learning is applied to the generation of 3D virtual architecture, particularly utilizing 3D Deep Generative Models (DGMs). In this context, two key scope notions warrant emphasis: Firstly, we underscore 3D DGMs due to their potential to efficiently generate designs aligned with diverse goals, thereby enriching content and affordances in the metaverse. We also include architectural research that generates 2D images with 3D transitions rather than stopping at 2D-style imaginary drawings. Secondly, our scope is confined to the generation task, distinguishing it from similar reconstruction and completion tasks, as generation involves actively creating 3D shapes using deep learning. This decision is informed by the prevalent application of generative tasks in architectural design, closely aligning with human cognition. Furthermore, in line with the virtual context of our inquiry, we prioritize articles where virtual environments can substitute physical ones rather than solely addressing real-world problems like Building Information Modeling (BIM).
@@ 2.1. Virtual Architecture and Design Discipline
2.1.1. Virtual Architecture: Mission and Scope.
@@
Our scrutiny reveals an overlapping scope between virtual architecture and human-building interaction (HBI), referring to a consistent goal between the social acceptability of metaverse (Lee et al., 2021b) and sustainability of HBI (Alavi et al., 2019). Specifically, HBI focuses on understanding and facilitating human interaction, adaptation, and emotional connection within the built environment (Alavi et al., 2019), whose sustainability highly echos the social acceptability of the metaverse. Furthermore, researchers have built a close correlation between these two notions relying on user perception in immersive virtual environments aiming at HBI goal (Heydarian et al., 2015). Additionally, according to conceptualizing HBI by Alavi et al., the built environment comprises social, spatial, and physical aspects (Alavi et al., 2019). This work also emphasizes that significant technologies driving HBI include ubiquitous computing, interactive architecture components, and sustainability. Notably, previous HBI surveys emphasize the inherent attributes of buildings, distinguishing them from traditional HCI (Alavi et al., 2019; Nembrini and Lalanne, 2017). Therefore, the HBI perspective offers a new approach to understanding the multifunctional nature of virtual buildings (Alavi et al., 2019), regarding encapsulating ubiquitous computing, human-centric interaction, and sustainability goals. We outline these essentials to define the mission and scope of virtual architecture in Table 1.
Table 1. Overview mission and scope of virtual architecture from five aspects, including spatial forms, user presence, socialization, interaction, and interoperability and scalability.
Discipline Description
Spatial forms
Incorporating diverse forms in terms of function, style, and purpose to ensure efficient production to achieve large-scale outcomes.
User presence, experience
Ensuring self-presence with a pleasant, understandable user experience, considering human perception (e.g., comfort, safety), behavioral demands (e.g., navigation, gathering activities), and environmental factors (e.g., lighting, sounds, atmosphere) in immersive environments.
Socialization, engagement
Supporting diverse and multiple social interactions, including activities, engagements, and goals for live, entertainment, and work. Providing collaboration opportunities with integrated tools for editing, altering, and customizing tasks.
Interactive elements
Designing flexible, dynamic, and real-time interactive human-centric buildings by utilizing ubiquitous data from environmental intelligence, such as responsive or interactive building elements.
Interoperability, scalability
Building interoperability and scalability across multiple platforms focusing on data security, user privacy, and resource management for sustainability.
@@ 2.1.2. Design Discipline
Figure 4. Virtual architecture projects with various forms. (a) Joris Putteneers, 2016. Synesthesia. The project creates a surreal and complex architectural construction by algorithmic simulation in Houdini  [^8] Source: Source:https://putteneersjoris.xyz/projects/synesthesia/synesthesia.html; (b) E-motion: a digital interface allows users to capture real-time data to interact with for a rethinking of co-living among various species  [^9] Fei Chen, Mochen Jiang, Haojun Cui, and Yuankai Wang, E-motion, 2020. Source: https://bproautumn2020.bartlettarchucl.com/rc18/e-motion; (c) Viviane Toraci Fiorella, Taza Celilia, Prandini Alvaro Campo. ”ISOS” in the “Volumetric Cinema” workshop by Current.CAM, 2022, which focuses on the interaction between virtual avatars and fantastically dramatic environments; (d) Tane Moleta and Mizuho Nishioka, the co-constructive project, ”Populating Virtual Worlds Together,” 2021, leading to an autonomous virtual space; (e) 3D volumetric architecture in virtual environments by utilizing BCI to capture affective-driven dynamic noise (Barsan-Pipu et al., 2020); (f) Current.CAM, VR gallery, 2021. Current. CAM’s VR exhibition is formed by continuous partitioned spaces with purely fluid blue, reinforcing the digital interface’s shaping on the human senses.
@@
    Building Form. First, deep generative models could form various formats, from solid mesh to the discrete point cloud, to support virtual architecture with flexible and inclusive representations. Generation and virtual techniques enable the construction and rendering of cutting-edge buildings in the virtual world by replying to proper datasets, such as Paper Architecture  [^10] Visionary architecture that could not be built in reality, only as drawings, collages, or models., and bionic architecture  [^11] Bionic architecture is usually computationally adapted to the structure or form of organic matter in nature. Design considerations for biomimetic architecture include living organisms’ physiological, behavioral, and structural adaptations., which previously not be implemented in reality. Second, leveraging deep generative models could form a diversified and rich shared online space since this technique not only incorporates the vast and various datasets contributed by users but also flexibly adjusts its framework or workflow. Space geometry has been expanded to space intelligence, representing the generative logic of forms transformatively. For instance, space form could be embedded in technical capabilities as assistance, such as algorithmic fluid form (Fig. 3(a)3(f)), spatial data with point cloud (Fig. 3(b)), interactive narrative (Fig. 3(c)), collaborative co-constructing (Fig. 3(d)), and creative voxel girds output (Fig. 3(e)).
@@
By highlighting these key points, we craft a virtual architecture guideline urging a broad understanding grounded in knowledge and expertise for virtual buildings in the metaverse.
@@ 2.2. Deep Generative Models and 3D Representations
Figure 5. The frameworks of GAN, VAE and Diffusion Models.
2.2.1. 3D Shape Generation.
Traditional deep generative models contribute to 3D shape generation alongside well-known models such as Generative Adversarial Networks (GANs) and variational autoencoders (VAEs). Other models like normalizing flows (N-Flows), recent diffusion probabilistic models (DDPMs), and energy-based models (EBMs) have also made significant contributions by learning from the similarity of given data. These deep generative models generate a tangible 3D object ready for rendering. It conveys a latent variable to a high-quality image. Although each model has benefits and has shown remarkable progress in recent years, the architecture domain primarily relies on GANs. In contrast, VAE and the latest diffusion models are designed for a few research. Considering the relevance, we thus introduce the GAN, VAE, and diffusion models in detail rather than an exhaustive list of models included in other CV survey articles (Fig.  5).
@@
2.2.2. 3D-Aware Image Synthesis.
This approaches extract latent vectors from the latent space and decodes them into a target representation by using GAN. Generally, the generation pipeline is for an image with 3D awareness as a result, and it also starts with an image as a generative source.
@@
Figure 6. The 3D representations for (a) Voxel grids, (b) Meshes, (c) Point cloud, (d) Neural fields.
@@
2.3. Generated Virtual Architecture by Deep Learning
@@
Reference Category Research Question 3D Representations Generative Models Key Focus
Dataset
Multi-Modality
Design Intuition
Generative Model
(Çakmak, 2022)
3D Solid
Extend design cognition
Point cloud/Mesh
GAN (with Autoencoder)
(Ondrej Vesel‘y, 2022)
3D Solid
Generate building massing
GAN
(Mueller, 2023)
3D Solid
Generate building massing
Point cloud/Voxel Grid
GAN
(Miguel et al., 2019)
3D Solid
Generation, manipulation and form finding of structural typologies
Wireframe
VAE
(Kahraman et al., 2021)
3D Solid
Solve design problems incorporating deep learning
Voxel grid
Base VAE & GAN
(Sebestyen et al., 2021)
3D Solid
New way to design models
Voxel grid
VAE
(Guida, 2023)
NLP-3D Solid
Multimodal generation with text input
Diffusion model
Table 2. This article elucidates four key focuses summarized from the research questions we reviewed for deep learning-assisted architectural design papers: datasets, multimodality, design intuition, and generative models (four columns on the right). A sample of the papers we reviewed is listed in the table as a reference for the process.
@@ 2.3.1. Related Works
Integrating deep learning into architectural design has significantly increased in recent years. Initially, neural networks in deep learning were widely utilized for style transfer in architectural exploration (depicted as 3D transposition in Fig. 7). For instance, most approaches employed style transfer to generate 2D images (e.g., floor plans or sections) by blending architectural styles with design intentions and manually assembling these images into 3D buildings as a post-processing step (Jaminet et al., 2022). However, this indirect method poses significant challenges for designers due to the absence of interpretative embeddings in each image, such as materiality, style, proportions, or program representation (Guida, 2023). Besides, a general lack of knowledge and expertise in generation techniques regarding frameworks and algorithms in the architectural field leads to underexploration in the design.
@@
Recent works have shifted focus from 2D images to 3D shapes, utilizing techniques such as 3D GAN, VAE, and 3D diffusion (depicted as 3D solid form generation in Fig. 7). Current approaches support architectural generation from different modality inputs, such as text, image, and 3D, leveraging different model features. Generally, the GAN utilizes labeled 3D datasets according to design meaning, such as architectural elements. VAE generates 3D models utilizing reduced 3D features recorded in latent space with an encoder-decoder. As for 3D diffusion, although its application in architectural design is very limited due to the rapid development of its multi-modal integration ability and a variety of methods, 3D diffusion has great application potential. Additionally, delving into 3D generation requires a higher technical threshold than simply applying generative models. For instance, in the generation process, GAN requires careful adjustment of hyperparameters for stability, network depth, width, and kernel size to ensure performance and output quality, as well as considerations of computational resources, dataset characteristics, and random noise input to the generator for training (Mueller, 2023).
@@
Both 2D and 3D methodologies shed light on the evolving trend towards designer-assisted AI workflows rather than mere AI-assisted design. Shi et al. underscored the importance of regulation and training by designers in such human-AI collaboration, alongside ideation, creation, visualization, and testing (Shi et al., 2023a). They outline four possibilities regarding regulation and training for better aligning with design goals: curating training datasets, providing semantic information, regulating AI’s roles and behaviors, and assisting AI in improving performance (Shi et al., 2023a). Architectural design experiments have also propelled such pioneering endeavors. For example, Huang proposed an experimental approach in semantic analytics through human interpretation (Huang et al., 2021), facilitating mutual learning between humans and AI to generate cultural and architectural meaning. Overall, this shift enables designers to contribute to future architecture in virtual and physical realms by transitioning from traditional design roles to ”technical design.”
@@
We identified four key focuses by thoroughly examining the literature on architectural design employing deep learning: dataset, multimodality, design intuition, and generative framework (see Table 2). Notably, our focus is specifically on architectural design and generation considering 3D form, aligning with the scope of our survey.
@@ 2.3.2. Dataset
Dataset preparation comprises data collection, data cleaning, and filtering, which is the first and foremost step in deep learning-assisted architectural design. However, 3D building datasets regarding specialized design and general applicability are currently scarce and underexplored. Despite being scrutinized, we only found three 3D building datasets (Ondrej Vesel‘y, 2022; Mueller, 2023; Peters et al., 2022). Selvaraju et al. first introduced BuildingNet, a large-scale collection of 3D building models with labeled exteriors (Selvaraju et al., 2021). Building upon this work, Mueller developed the HouseNet dataset, a selective collection of preprocessed models, which was subsequently utilized in her design work (Mueller, 2023). Besides, Peters et al. introduced the 3DBAG dataset, containing 3D building models of the Netherlands (Peters et al., 2022), featuring multiple levels of detail. The 3DBAG dataset was further leveraged by Vesely in deep learning-assisted architectural design utilizing GAN for generating 3D geometry building (Ondrej Vesel‘y, 2022), showcasing its potential in design.
Therefore, architectural design research has to prioritize building 3D datasets through manual pre-processing before the generation process. This includes generating models using various tools, such as generating massive models approach (Sebestyen et al., 2021), scanning with specialized scanners (Çakmak, 2022), segmenting architectural elements (Çakmak, 2022; Ondrej Vesel‘y, 2022), and formatting and transforming data (Mueller, 2023). The lack of adequate and unified datasets hampers further exploration of broad user affordances and democratized tool development.
@@ 2.3.3. Multimodality
Figure 7. Various architecture design methods by deep learning towards multimodality, then divided into two groups, namely 3D transposition and 3D solid form. 3D transposition leverages 2D generative images to transpose into 3D form in an additional manual or computational way. In contrast, 3D solid form generation leverages 3D deep generative models, including 3D GANs, VAE, and 3D diffusion, to generate 3D form directly.
Multimodality leverages natural human communication capabilities by integrating multiple modalities, such as speech, gestures, and visuals, within a single channel. The current literature underscores the significance of multimodality in deep learning methods, particularly in utilizing human understanding as input, such as descriptive prompts and stylized images. Generally, these diverse design methods can be classified as follows, which can also be divided into two groups: 2D and 3D generation approaches with deep learning (3D transposition and 3D solid generation as shown in Fig. 7):
@@ 2.3.4. Design Intuition
Design intuition refers to processing generative techniques to cooperate with intended design goals (Koh, 2022; Çakmak, 2022), determining the architectural meaning of design outputs through extended interpretability. Although many factors can affect the design results aiming to enhance design intuition, we list the two most mainstream ones: Generative images could be tailored to the design purpose in 3D transposition approaches by incorporating the desired style. For instance, Ren et al. utilized StyleGAN to blend the Gothic architectural style with an original plan by inputting the real essence of a Gothic-style floor plan (Ren and Zheng, 2020).
Preprocessing datasets could also facilitate the design intuition in both 3D transposition and 3D solid form approaches (Koh, 2022; Çakmak, 2022), which ensures the dataset contains the necessary and accurate architectural elements imbued with ”design meaning.” However, the current approach often involves preprocessing data by designers according to specific design requirements, demanding significant human and computational effort to drive interpretability toward design objectives. For example, Çakmak conducted architectural element segmentation on scanned point clouds to generate meaningful components with GAN and an encoder-decoder architecture (Çakmak, 2022).

### v4

@@ 1.3. Scope and Structure
While the intersection of deep learning and architecture encompasses various aspects, including design cognition (e.g., feature extraction, image recognition, design evaluation), design generation (e.g., floor plans, rendering, form generation), and assisting tools (e.g., efficiency optimization, strategy prediction), our investigation focuses specifically on articles where deep learning is applied to the generation of 3D virtual architecture, particularly utilizing 3D Deep Generative Models (DGMs). In this context, two key scope notions warrant emphasis: Firstly, we underscore 3D DGMs due to their potential to efficiently generate designs aligned with diverse goals, thereby enriching content and affordances in the virtual environment. We also include architectural research that generates 2D images with 3D transitions rather than stopping at 2D-style imaginary drawings. Secondly, our scope is confined to the generation task, distinguishing it from similar reconstruction and completion tasks, as generation involves actively creating 3D shapes using deep learning. This decision is informed by the prevalent application of generative tasks in architectural design, closely aligning with human cognition. Furthermore, in line with the virtual context of our inquiry, we prioritize articles where virtual environments can substitute physical ones rather than solely addressing real-world problems like Building Information Modeling (BIM).
@@ 2.1. Virtual Architecture and Design Discipline
2.1.1. Virtual Architecture: Mission and Scope
@@
Our scrutiny reveals an overlapping scope between virtual architecture and human-building interaction (HBI), referring to a consistent goal between the social acceptability of metaverse (Lee et al., 2021b) and sustainability of HBI (Alavi et al., 2019). Specifically, HBI focuses on understanding and facilitating human interaction, adaptation, and emotional connection within the built environment (Alavi et al., 2019), whose sustainability highly echos the social acceptability of the metaverse. Furthermore, researchers have built a close correlation between these two notions relying on user perception in immersive virtual environments aiming at HBI goal (Heydarian et al., 2015). Additionally, according to conceptualizing HBI by Alavi et al., the built environment comprises social, spatial, and physical aspects (Alavi et al., 2019). This work also emphasizes that significant technologies driving HBI include ubiquitous computing, interactive architecture components, and sustainability. Notably, previous HBI surveys emphasize the inherent attributes of buildings, distinguishing them from traditional HCI (Alavi et al., 2019; Nembrini and Lalanne, 2017). Therefore, the HBI perspective offers a new approach to understanding the multifunctional nature of virtual buildings (Alavi et al., 2019), regarding encapsulating ubiquitous computing, human-centric interaction, and sustainability goals. We outline these essentials to define the mission and scope of virtual architecture in Table 1 (Appendix).
@@ 2.1.2. Design Discipline
Figure 4. Virtual architecture projects exhibit various forms: (a) Joris Putteneers’ ”Synesthesia”  [^9] Source: https://putteneersjoris.xyz/projects/synesthesia/synesthesia.html(2016) creates a surreal and complex architectural construction through algorithmic simulation in Houdini; (b) ”E-motion”  [^10] Source: https://bproautumn2020.bartlettarchucl.com/rc18/e-motion(2020) by Fei Chen et al. features a digital interface allowing real-time data interaction for rethinking co-living among species; (c) ”ISOS” by Viviane Toraci Fiorella, Taza Celilia, and Prandini Alvaro Campo in the ”Volumetric Cinema” workshop by Current.CAM (2022) explores the interaction between virtual avatars and dramatic environments; (d) Tane Moleta and Mizuho Nishioka’s collaborative project (2021), ”Populating Virtual Worlds Together,” leads to an autonomous virtual space; (e) Utilizing BCI to capture affective-driven dynamic noise, Barsan et al. (2020) demonstrate 3D volumetric architecture in virtual environments (Barsan-Pipu et al., 2020); (f) Current.CAM’s VR gallery (2021) features continuous partitioned spaces with fluid blue, reinforcing the shaping of digital interfaces on the human senses.
@@
    Building Form. First, deep generative models could form various formats, from solid mesh to the discrete point cloud, to support virtual architecture with flexible and inclusive representations. Generation and virtual techniques enable the construction and rendering of cutting-edge buildings in the virtual world by replying to proper datasets, such as Paper Architecture  [^11] Visionary architecture that could not be built in reality, only as drawings, collages, or models., and bionic architecture  [^12] Bionic architecture is usually computationally adapted to the structure or form of organic matter in nature. Design considerations for biomimetic architecture include living organisms’ physiological, behavioral, and structural adaptations., which previously not be implemented in reality. Second, leveraging deep generative models could form a diversified and rich shared online space since this technique not only incorporates the vast and various datasets contributed by users but also flexibly adjusts its framework or workflow. Space geometry has been expanded to space intelligence, representing the generative logic of forms transformatively. For instance, space form could be embedded in technical capabilities as assistance, such as algorithmic fluid form (Fig. 4(a)4(f)), spatial data with point cloud (Fig. 4(b)), interactive narrative (Fig. 4(c)), collaborative co-constructing (Fig. 4(d)), and creative voxel girds output (Fig. 4(e)).
@@
By highlighting these key points, we craft a virtual architecture guideline urging a broad understanding grounded in knowledge and expertise for virtual buildings.
@@ 2.2. Deep Generative Models and 3D Representations
Figure 5. The frameworks of GANs, VAEs, and Diffusion Models.
2.2.1. 3D Shape Generation
Traditional deep generative models contribute to 3D shape generation alongside well-known models such as Generative Adversarial Networks (GANs) and variational autoencoders (VAEs). Other models like normalizing flows (N-Flows), recent diffusion probabilistic models (DDPMs), and energy-based models (EBMs) have also made significant contributions by learning from the similarity of given data. These deep generative models generate a tangible 3D object ready for rendering. It conveys a latent variable to a high-quality image. Although each model has benefits and has shown remarkable progress in recent years, the architecture domain primarily relies on GANs. In contrast, VAEs and the latest diffusion models are designed for a few research. Considering the relevance, we thus introduce the GANs, VAEs, and diffusion models in detail rather than an exhaustive list of models included in other CV survey articles (Fig.  5).
@@
2.2.2. 3D-Aware Image Synthesis
This approach extracts latent vectors from the latent space and decodes them into a target representation by using GAN. Generally, the generation pipeline is for an image with 3D awareness as a result, and it also starts with an image as a generative source.
@@
Figure 6. The figure illustrates various 3D representations commonly used in generated architecture: (a) Voxel grids, (b) Meshes, (c) Point clouds, and (d) Neural fields.
@@
2.3. Generated Virtual Architecture with Deep Learning
@@ 2.3.1. Related Works
Integrating deep learning into architectural design has significantly increased in recent years. Initially, neural networks in deep learning were widely utilized for style transfer in 2D drawings of architectural design (depicted as 3D transposition in Fig. 7), such as floor plans (Chaillou, 2022, 2020), sections (Yu, 2020), and facades (Mohammad et al., 2019; Dong et al., 2023). Most approaches employ a zero-sum game of these neural networks to blend architectural styles with design intentions and then manually assemble these images into 3D buildings as a post-processing step (Jaminet et al., 2022). However, this indirect method poses significant challenges for designers due to the absence of interpretative embeddings in each image, such as materiality, style, proportions, or program representation (Guida, 2023). Additionally, a general lack of knowledge and expertise in generation techniques regarding frameworks and algorithms in the architectural field leads to underexploration in the design.
@@
Recent works have shifted focus from 2D images to 3D shapes, utilizing techniques such as 3D GANs, VAEs, and 3D diffusion models (depicted as 3D solid form generation in Fig. 7). Current approaches support architectural generation from different modality inputs, such as text, image, and 3D, leveraging different model features. Generally, GAN utilize labeled 3D datasets according to design meaning, such as architectural elements. VAEs generate 3D models utilizing reduced 3D features recorded in latent space with an encoder-decoder. Although the application of 3D diffusion models in architectural design is limited, their multi-modal integration capability and diverse methods suggest significant potential for future use. Additionally, delving into 3D generation requires a higher technical threshold than simply applying generative models. For instance, in the generation process, GANs require careful adjustment of hyperparameters for stability, network depth, width, and kernel size to ensure performance and output quality. Additionally, considerations of computational resources, dataset characteristics, and random noise input to the generator are essential for effective training (Mueller, 2023).
@@
Both 2D and 3D methodologies shed light on the evolving trend towards designer-assisted AI workflows rather than mere AI-assisted design. Shi et al. underscored the importance of regulation and training by designers in such designer-AI collaboration, alongside ideation, creation, visualization, and testing (Shi et al., 2023a). They outline four possibilities regarding regulation and training for better aligning with design goals: curating training datasets, providing semantic information, regulating AI’s roles and behaviors, and assisting AI in improving performance (Shi et al., 2023a). Architectural design experiments have also propelled such pioneering endeavors. For example, Huang proposed an experimental approach in semantic analytics through human interpretation (Huang et al., 2021), facilitating mutual learning between humans and AI to generate cultural and architectural meaning. Overall, this shift enables designers to contribute to future architecture in virtual and physical realms by transitioning from traditional design roles to ”technical design.”
@@
We identified four key focuses by thoroughly examining the literature on architectural design employing deep learning: dataset, multimodality, design intuition, and generative framework (see Table 2, Appendix). Notably, our focus is specifically on architectural design and generation considering 3D form, aligning with the scope of our survey.
@@ 2.3.2. Dataset
Dataset preparation comprises data collection, data cleaning, and filtering, which is the first and foremost step in deep learning-assisted architectural design. However, 3D building datasets regarding specialized design and general applicability are currently scarce and underexplored. Despite being scrutinized, we found only three building datasets that can aid in 3D form generation (Ondrej Vesel‘y, 2022; Mueller, 2023; Peters et al., 2022). Selvaraju et al. first introduced BuildingNet, a large-scale collection of 3D building models with labeled exteriors (Selvaraju et al., 2021). Building upon this work, Mueller developed the HouseNet dataset, a selective collection of preprocessed models, which was subsequently utilized in her design work (Mueller, 2023). Besides, Peters et al. introduced the 3DBAG dataset, containing 3D building models of the Netherlands (Peters et al., 2022), featuring multiple levels of detail. The 3DBAG dataset was further leveraged by Vesely in deep learning-assisted architectural design utilizing heatmap data for generating 3D geometry building (Ondrej Vesel‘y, 2022), showcasing its potential in design.
Therefore, architectural design research has to prioritize building 3D datasets through manual pre-processing before the generation process. This includes generating models using various tools, such as generating massive models approach (Sebestyen et al., 2021), scanning with specialized scanners (Çakmak, 2022), segmenting architectural elements (Çakmak, 2022), and formatting and transforming data (Mueller, 2023). The lack of adequate and unified datasets hampers further exploration of broad user affordances and democratized tool development.
@@ 2.3.3. Multimodality
Figure 7. 3D transposition leverages 2D generative images to transpose into 3D form in an additional manual or computational way. In contrast, 3D solid form generation leverages 3D deep generative models, including 3D GANs, VAEs, and 3D diffusions, to generate 3D form directly.
Multimodality leverages natural human communication capabilities by integrating multiple modalities, such as speech, gestures, and visuals, within a single channel. The current literature underscores the significance of multimodality in deep learning methods, particularly in utilizing human understanding as input, such as descriptive prompts and stylized images. Generally, these diverse design methods can be classified as follows, which can also be divided into two groups: 2D and 3D generative approaches with deep learning (3D transposition and 3D solid generation as shown in Fig. 7):
@@ 2.3.4. Design Intuition
Design intuition refers to processing generative techniques to cooperate with intended design goals (Koh, 2022; Çakmak, 2022), determining the architectural meaning of design outputs through extended interpretability. Although many factors can affect the design results aiming to enhance design intuition, we list the two most mainstream ones: Generative images could be tailored to the design purpose in 3D transposition approaches by incorporating the desired style. For instance, Ren et al. utilized a StyleGAN to blend the Gothic architectural style with an original plan by inputting the real essence of a Gothic-style floor plan (Ren and Zheng, 2020).
Preprocessing datasets could also facilitate the design intuition in both 3D transposition and 3D solid form approaches (Koh, 2022; Çakmak, 2022), which ensures the dataset contains the necessary and accurate architectural elements imbued with ”design meaning.” However, the current approach often involves preprocessing data by designers according to specific design requirements, demanding significant human and computational effort to drive interpretability toward design objectives. For example, Çakmak conducted architectural element segmentation on scanned point clouds to generate meaningful components with a GAN and an encoder-decoder architecture (Çakmak, 2022).